Determine all the points on the curve y = 2x3 + 3x2 - 18x + 3 where the slope of the tangent line is -6.


1
Expert's answer
2021-03-31T13:54:09-0400

Consider the function "y=2x^3+3x^2-18x+3"


Differentiate with respect to "x" as,


"\frac{dy}{dx}=\frac{d}{dx}(2x^3+3x^2-18x+3)"


"=2(3x^2)+3(2x)-18(1)+0"


"=6x^2+6x-18"


Set "\frac{dy}{dx}=-6" (which is the slope of the tangent line) and solve for "x" as,


"6x^2+6x-18=-6"


"x^2+x-3=-1"


"x^2+x-2=0"


"x^2+2x-x-2=0"


"x(x+2)-1(x+2)=0"


"(x-1)(x+2)=0"


"x=-2,1"


At "x=-2,y=35" and at "x=1,y=-10"


Therefore, the points on the curve are: "(-2,35),(1,-10)".